This scatterplot shows data on a sample of 20 toddlers observed over several months at a nursery school. Can you predict how many calories a child consumes based on how long they remain at the lunch table? “Time” is the average number of minutes a child spent at a table when lunch was served. “Calories” is the average number of calories consumed by the child during lunch. The least squares regression equation is: calories = 560.65125 – 3.0770733 × time. 1. What is the predicted number of calories consumed for a child who spends 40 minutes at the table? 437.6 calories 123.1 calories 548.3 calories
Answer #1

Calories =560.65125 - 3.0770733Time

When time = 40 min

Predicted calories = 560.65125 - 3.0770733*40 = 437.6

Answer: 437.6 calories